OpenStudy (anonymous):

What is the discriminant of f(x) = -2x^2 +3x +5 @Aleah.wood1398

OpenStudy (anonymous):

one second!

OpenStudy (campbell_st):

for a quadratic \[ax^2 + bx + c\] the discriminant is \[\Delta = b^2 - 4ac\] you have a = -2, b = 3 and c = 5 sibstitute and calculate

OpenStudy (anonymous):

Discriminant=b^2-4(a)(c)=3^2-4(-2)(5)=49

OpenStudy (anonymous):

A rectangle shed covers 800 square feet of ground. it is 10 ft longer than it is wide. What is the width of the shed to the nearest tenth of a foot? @Aleah.wood1398

OpenStudy (anonymous):

this is a tricky question

OpenStudy (anonymous):

It really is

OpenStudy (campbell_st):

start by drawing a diagram  include the given information and get an equation for the area. so what can you say about the area in terms of the width, w

OpenStudy (anonymous):

the width is multiplied twice in the area?

OpenStudy (anonymous):

that makes it alot easier^^

OpenStudy (campbell_st):

can you distribute \[(w + 10) \times w = ?\] and you will need the general quadratic formula for the solution

OpenStudy (anonymous):

which is what?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

can you expand or distribute ?

OpenStudy (anonymous):

distribute

OpenStudy (campbell_st):

it means to multiply the term outside by both terms inside the brackets

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Im still not sure how to solve it

OpenStudy (campbell_st):

so what did you get...?

OpenStudy (anonymous):

do i do (800 +10) x800?

OpenStudy (campbell_st):

using the given area information you get \[(w + 10) \times w = 800\] you still have the same problem... you need to distribute

OpenStudy (anonymous):

subtract 10 from both sides?

OpenStudy (campbell_st):

no....the process of expanding or distributing involves multiplying each term inside the brackets by the term outside the brackets so what is \[w \times w = ?~~~~ and ~~~~~ 10 \times w = ?\]

OpenStudy (anonymous):

w^2 and 10 w

OpenStudy (campbell_st):

great... so now \[w^2 + 10w = 800\] rewriting the equation \[w^2 + 10w - 800 = 0\] now solve for w... you'll need the general quadratic formula

OpenStudy (anonymous):

what is the general formula?

OpenStudy (campbell_st):

well its \[w = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] for a quadratic \[aw^2 + bw + c = 0\] so match the information, substitute and then calculate an answer. You only need the positive solution as measurements can't be negative

OpenStudy (anonymous):

ill be using the quadratic right?

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

i got aw^2 +10w-800=0

OpenStudy (campbell_st):

why aw^2... I gave you the equation... several posts ago its \[w^2 + 10w - 800 = 0\] so when you compare it to the general form \[aw^2 + bw + c = 0\] what are the values of a, b and c

OpenStudy (anonymous):

i thought I ws to stay in the equation so its just w^2 +10w-800=0